# L2M-Reg: Building-level Uncertainty-aware Registration of Outdoor LiDAR Point Clouds and Semantic 3D City Models

### v2

@@ 3.2 Reliable Plane Correspondence Establishment
After data preprocessing, accurately extracting a suited LiDAR plane segment $L_{i}$ from neighboring point cloud $N_{i}$ to construct reliable correspondences remains a challenge. To address this, an automated solution is developed, comprising two main parts: representative region localization in Section 3.2.1 and plane segment extraction in Section 3.2.2.
@@ 3.2.1 Automated Representative Region Localization
    Detecting the largest plane from $N_{i}$ based on Random Sample Consensus (RANSAC) as shown in red part in Fig 4, denoted as $P_{\text{max}}$. The distance threshold $T_{{dis}}$ used in the RANSAC is determined based on the registration residual $e$ of the entire input LiDAR point clouds, defined as $T_{{dis}}=e$. The residual $e$ can be derived from the performance of the multi-station point cloud registration.
@@
    The $Z$ coordinates of all the points in $P_{\text{max}}$ are then sorted in ascending order, and the value at the 10th and 90th percentiles is selected as the lower and upper height bounds as shown in solid and dashed black lines in Fig 4, respectively, denoted as $Z_{\min}$ and $Z_{\max}$.
@@
    The angle $\alpha$ between the normal vector of $P_{\text{max}}$ and the ground plane is then calculated. If $\alpha$ exceeds angle threshold $T_{{\alpha}}$, $P_{\text{max}}$ is considered part of the exterior facade. In this case, all points in $N_{i}$ whose $Z$ coordinates are larger than $Z_{\min}$ are removed, and the $P_{\text{max}}$ extraction process is also recognized as valid. The remaining point cloud is then taken as the updated $N_{i}$ for the next iteration as shown in Fig 4.
@@ 3.2.2 Automated Plane Segment Extraction
More details are shown in Algorithm 2. The proposed plane extraction algorithm is better suited than the classic RANSAC to extract representative planes from irregular and rough areas in real-world scenarios. Typically, RANSAC identifies a set of inliers that satisfy predefined threshold. It outputs the plane corresponding to the largest inlier set while discarding all outliers by default.
@@
However, in practice, the building surface may contain many small-scale irregularities or exhibit pronounced roughness. In addition, due to registration residual introduced by multi-station scanning, the input point clouds used for plane extraction may suffer from local layer separation. In such cases, some of the points rejected by RANSAC as outliers may, in fact, still be suited and usable. By leveraging GC check as shown in Fig 5, the proposed algorithm supplements the inlier set with structurally coherent points from other sets and outliers. This enhances the representativeness of the final extracted plane while preserving the overall consistency of extraction quality.
@@
After the above process is completed, the output is the final LiDAR plane segment $L_{i}$, whose corresponding model-derived plane $M_{i}$ has already been fixed during preprocessing. However, it is important to note that all current plane correspondences only cover building exterior facade or plinth area. To achieve the 6 DoF parameter estimation, additional correspondence from the ground area is still required. This issue will be addressed in the following Section 3.3.
3.3 Pseudo-plane-constraint Gauss–Helmert Model
@@
However, in practice, the effectiveness of DTM data is often limited. Most DTM data are derived from Airborne Laser Scanning (ALS), which provides significantly lower resolution and geometric accuracy than the TLS or MLS data used for facade-based correspondences (Macay Moreia et al., 2013), as shown in Figure 6. As a result, the relatively low vertical accuracy of DTM-derived correspondences will introduce uncertainty into the overall transformation estimation. In particular, errors in the vertical direction may propagate and adversely affect the estimation of other components, such as horizontal translation and rotation, ultimately compromising the accuracy and stability of the full parameter estimation. This is primarily due to the coupled nature of estimation using plane-based Gauss–Helmert Model (GHM), where vertical inaccuracies can distort the orientation of fitted planes and bias the least-squares optimization process, thereby contaminating the estimation of parameters in other directions.
To address this, a decoupled 2D–3D estimation strategy is proposed to effectively prevent vertical inaccuracies from propagating into other components. Specifically, a pseudo-plane correspondence is introduced to replace the unreliable correspondence from DTM and provide additional constraint in the GHM adjustment.
@@
Figure 6: Introduced pseudo-plane correspondence. $L_{i}$ and $M_{i}$ denote the LiDAR-derived and model-derived planes, respectively, and serve as facade-based correspondences for registration. The dashed lines represent the introduced pseudo-plane correspondence, which serve as substitutes for the original constraints derived from the LiDAR ground points and model ground points (DTM), indicated by red and blue round dots, respectively.
@@
$\mathbf{A}_{[n\times 7]}=\left.\frac{\partial f}{\partial\mathbf{X}}\right|_{\mathbf{X}_{0},\mathbf{V}_{0}}\qquad\mathbf{B}_{[n\times 3n]}=\left.\frac{\partial f}{\partial\mathbf{V}}\right|_{\mathbf{X}_{0},\mathbf{V}_{0}}\qquad\mathbf{C}_{[1\times 7]}=\left.\frac{\partial c}{\partial\mathbf{X}}\right|_{\mathbf{X}_{0}}$ (11)
@@
The primary advantage of this strategy lies in its ability to perform reliable parameter estimation even in the absence of real ground-plane correspondences. More importantly, it effectively decouples the influence of low-quality observations—such as those derived from DTM data—from high-quality observations, preventing error propagation across components and maximizing the use of reliable facade-based correspondences.
On the one hand, such design remains adaptable to ground data of varying types, accuracies, and acquisition sources. This adaptability becomes even more evident in scenarios with sloped or uneven ground surfaces, where simply constraining or omitting the parameter in a given direction—such as $t_{z}$ in the GHM—may still introduce additional errors. On the other hand, rather than omitting the estimation of $t_{z}$ and solving only for five parameters, the pseudo-plane strategy offers greater flexibility and adaptability, making it readily extendable to other facade orientations. Overall, the pseudo-plane strategy delivers a balanced design that preserves estimation robustness while maintaining adaptability to diverse scenarios.
@@ 3.4 Adaptable Vertical Translation Estimation
Figure 7: Vertical translation estimation
In this way, the vertical translation $t_{z}$ can be easily estimated from DTM data while preserving a high degree of flexibility, benefiting from the aforementioned 2D–3D decoupled estimation strategy. For instance, if more accurate vertical information becomes available—such as GNSS observations, total station measurements, or other high-fidelity ground models—the DTM can be seamlessly substituted using the same method. This design greatly enhances the adaptability of the overall framework to heterogeneous data sources and ensures robustness under varying data availability conditions.
@@ 4.1 Data Preparation
To validate and evaluate the proposed L2M-Reg, three datasets (as shown in Figure 8), named as TUM0501 Building, Pinakothek, and Street Building, are collected, comprising LiDAR point clouds of buildings based on TLS or MLS from the urban areas of Munich. Besides, the corresponding publicly accessible LoD2 models are obtained from the Bavarian Surveying Administration [^2] https://geodaten.bayern.de/opengeodata/. These datasets vary in point cloud size, building size, and data acquisition methods.
Figure 8: Overview of three datasets. The upper part shows the LoD2 semantic models, with planes colored for visual distinction. The lower part displays the corresponding point clouds, colored by intensity.
In particular, due to the absence of ground data in the publicly accessible LoD2 models, ground information for the TUM0501 Building and Pinakothek datasets which are part of the TUM2TWIN [^3] https://tum2t.win/ (Wysocki et al., 2025) is supplemented by publicly available DTM data with a grid width of 1 m from the Bavarian Surveying Administration. The Street Building dataset utilizes an additional georeferenced road model. Further details of the datasets are summarized in Table 1.
Table 1: Basic Information on Three Datasets
TUM0501 Building Pinakothek Street Building
Number of Points 2,272,810 10,367,229 5,000,509
Scanning Area About 1,200 m2 About 36,000 m2 About 1,500 m2
Scanning System Leica ScanStation P50 Leica ScanStation P50 Z+F FlexScan 22
Scanning Technique TLS TLS MLS
Ground Data Source DTM DTM Road Model
@@ 4.2 Experiment Design
In addition to the proposed L2M-Reg, four existing methods — GICP (Segal et al., 2009), TriICP (Chetverikov et al., 2002), PLADE (Chen et al., 2019) and Scantra (Wujanz et al., 2018) — are also applied to the three datasets for performance evaluation and comparison. GICP and TriICP are chosen as leading robust ICP-variants, known for their reliability in fine-registration tasks. PLADE and Scantra are included as the two leading plane-based registration methods. Since these registration methods are inherently designed for point clouds rather than parametric models, the LoD2 models from the three datasets were converted into point clouds with uniform density to facilitate registration using the four methods mentioned above.
L2M-Reg was implemented in C++ using Point Cloud Library (PCL) (Rusu and Cousins, 2011). Standard implementations of GICP and TriICP in PCL (version 1.13.0) were utilized, while PLADE was implemented based on its open-source release [^4] https://github.com/chsl/PLADE (Chen et al., 2019). Scantra was tested on the basis of the latest version 3.4 released in 2025555https://www.technet-gmbh.com/en/products/scantra/scantra-news/scantra-release-34-static-kinematic-and-polar-are-ready-for-download/. All methods were executed under identical hardware conditions (Intel Xeon(R) W-2223 CPU@3.60GHz with 4 Cores and 64 GB of RAM).
@@ 4.3 Evaluation Metrics
where $n_{H}$ and $n_{V}$ denote the number of check points used for assessing horizontal and vertical error, respectively. For each check point $p$, $\delta_{\mathrm{M3C2}}(p)$ represents the M3C2 distance between $p$ and the reference model. Similarly, based on Equation 14 and Equation 15, the standard deviation of M3C2 distances in the horizontal and vertical directions can be computed by
@@ 4.4 Comparison Results
It is important to note that, unlike conventional evaluations based on the overall geometric closeness of the entire building, this study — motivated by the focus on model uncertainty — adopts the building footprint area as the primary basis for evaluation. This choice stems from the fact that, during LoD2 model generation, the near-ground footprint area is typically the most suited and accurately reflects the building’s true location, especially when there is a horizontal offset between plinth and facade. Accordingly, all subsequent evaluations in this study assess horizontal accuracy merely based on the distance between the aligned LiDAR point clouds and LoD2 models in the footprint area.
@@ 4.4.1 Qualitative Comparison
Figures 9, 10 and 11 present the valid registration results in three datasets. During testing, TriICP, GICP, and L2M-Reg consistently produced valid results, whereas PLADE failed to do so for the TUM0501 Building dataset. Scantra also exhibited suboptimal performance across three datasets. Possible reasons for these shortcomings are discussed in Section 5.
@@
As shown in Figure 9, the TUM0501 Building dataset yields valid registration results using TriICP, GICP, and L2M-Reg. The results of PLADE and Scantra are not included, as neither produces valid results in this case. In the horizontal direction, particularly within the building footprint areas outlined by dashed circles, L2M-Reg achieves the highest consistency compared to TriICP and GICP. In the vertical direction, it also shows superior alignment with the ground model, further highlighting its advantage in registration accuracy.
Figure 9: Registration performance of different methods in the TUM0501 dataset
@@
Figure 10 shows all valid registration results for the Pinakothek dataset. In the horizontal direction, L2M-Reg consistently demonstrates higher accuracy within the footprint area, as illustrated in Patch 1 to Patch 5 (outlined by dashed circles). Notably, in Patch 6, L2M-Reg exhibits a visible deviation from the reference model compared to the two ICP-based methods. This discrepancy is primarily attributed to scale inaccuracies in the LoD2 model, resulting in dimensional mismatches between the model and the point cloud. Despite this, L2M-Reg still achieves the globally optimal registration, as evidenced by Patch 1 being parallel to Patch 6, where L2M-Reg again provides a better global fit. In the vertical direction, as shown in Patch 7 to Patch 9, L2M-Reg demonstrates the best alignment between point clouds and DTM.
Figure 10: Registration performance of different methods in the Pinakothek Dataset
@@
In the Street Building dataset, shown in Figure 11, L2M-Reg consistently delivers the most accurate alignment in the horizontal direction, particularly in Patch 1 and Patch 4. The similar slight dimensional mismatches between the models and the point clouds are observed in Patch 2 and Patch 3, as seen in the Pinakothek dataset. These mismatches are again attributed to scale inaccuracies introduced in the original models. Nevertheless, L2M-Reg achieves a globally optimal result, with the registered point clouds (in red) centrally aligned between the two corresponding facades of the LoD2 models in Patch 2 and Patch 3.
In the vertical direction (e.g., Patch 6), L2M-Reg achieves the best alignment with the reference model in most regions. However, in certain areas like Patch 5, it does not produce optimal results. This results from the road being originally described according to the OpenDRIVE standard (ASAM, 2024; Kutsch et al., 2024), which models individual lanes with overlapping surface geometries and continuous curvature. The parametric surface representations are then sampled to explicit planar surface geometries as part of the conversion to CityGML 3.0 (Schwab et al., 2020). Although unintended overlapping surface geometries introduce minor instability in the evaluation of vertical accuracy, this demonstrates that L2M-Reg can also directly incorporate further object classes of the model, such as the road surface, into the vertical alignment process. Since CityGML 3.0 enables a comprehensive and redundancy-free representation of the road space, such ambiguities are not introduced in native mapping processes.
@@
In general, the qualitative comparison results indicate that L2M-Reg outperforms the other methods in both horizontal and vertical consistency in the three datasets. More detailed quantitative evaluation results are presented in Section 4.4.2.

### v3

@@ 3.2 Reliable Plane Correspondence Establishment
After data preprocessing, accurately extracting a suitable LiDAR plane segment $L_{i}$ from neighboring point cloud $N_{i}$ to construct reliable correspondences remains a challenge. To address this, an automated solution is developed, comprising two main parts: representative region localization in Section 3.2.1 and plane segment extraction in Section 3.2.2. In general, the input of this step consists of the point clouds covering the entire facade region, including parts of the overhanging roof, the exterior facade, and, most importantly, the wall plinth, as shown in Figure 1. The output is the extracted plane segment $L_{i}$ (wall plinth area). Since the wall plinth is the primary source of the model uncertainty addressed in this study.
@@ 3.2.1 Automated Representative Region Localization
    Detecting the largest plane from $N_{i}$ based on Random Sample Consensus (RANSAC) as shown in the red part in Figure 4, denoted as $P_{\text{max}}$. The distance threshold $T_{{dis}}$ used in the RANSAC is determined based on the registration residual $e$ of the entire input LiDAR point clouds, defined as $T_{{dis}}=e$. The residual $e$ can be derived from the performance of the multi-station point cloud registration.
@@
    The $Z$ coordinates of all the points in $P_{\text{max}}$ are then sorted in ascending order, and the value at the 10th and 90th percentiles is selected as the lower and upper height bounds as shown in solid and dashed black lines in Figure 4, respectively, denoted as $Z_{\min}$ and $Z_{\max}$.
@@
    The angle $\alpha$ between the normal vector of $P_{\text{max}}$ and the ground plane is then calculated. If $\alpha$ exceeds angle threshold $T_{{\alpha}}$, $P_{\text{max}}$ is considered part of the exterior facade. In this case, all points in $N_{i}$ whose $Z$ coordinates are larger than $Z_{\min}$ are removed, and the $P_{\text{max}}$ extraction process is also recognized as valid. The remaining point cloud is then taken as the updated $N_{i}$ for the next iteration, as shown in Figure 4.
@@ 3.2.2 Automated Plane Segment Extraction
More details are shown in Algorithm 2. The proposed plane extraction algorithm is better suited than the classic RANSAC to extract representative planes from irregular and rough areas in real-world scenarios.
@@
Typically, RANSAC identifies a set of inliers that satisfy a predefined threshold. It outputs the plane corresponding to the largest inlier set while discarding all outliers by default. However, in practice, the building surface may contain many small-scale irregularities or exhibit non-planar components. In addition, due to registration residual introduced by multi-station scanning, the input point clouds used for plane extraction may suffer from local layer separation. In such cases, simply using RANSAC for plane extraction yields suboptimal results, producing multiple discrete planes. Some of the points rejected by RANSAC as outliers may, in fact, still be suited and usable. By leveraging GC check as shown in Figure 5, the proposed algorithm supplements the inlier set with structurally coherent points from other sets and outliers. More specifically, the proposed plane extraction algorithm extends RANSAC with similarity-based clustering and merging steps. The GC check is then applied to further improve the reliability of the extracted planes. As a result, the method remains applicable when exterior facades contain partially non-planar components.
@@
After the above process is completed, the output is the final LiDAR plane segment $L_{i}$, whose corresponding model-derived plane $M_{i}$ has already been fixed during preprocessing. However, it is important to note that all current plane correspondences only cover the building exterior facade or plinth area. To achieve the 6 DoF parameter estimation, additional correspondence from the ground area is still required. This issue will be addressed in the following Section 3.3.
3.3 Pseudo-plane-constraint Gauss-Helmert Model
@@
However, in practice, the effectiveness of DTM data is often limited. Most DTM data are derived from Airborne Laser Scanning (ALS), which provides significantly lower resolution and geometric accuracy than the TLS or MLS data used for facade-based correspondences (Macay Moreia et al., 2013), as shown in Figure 6. As a result, the relatively low vertical accuracy of DTM-derived correspondences will introduce uncertainty into the overall transformation estimation. In particular, errors in the vertical direction may propagate and adversely affect the estimation of other components, such as horizontal translation and rotation, ultimately compromising the accuracy and stability of the full parameter estimation. This is primarily due to the coupled nature of estimation using plane-based Gauss-Helmert Model (GHM), where vertical inaccuracies can distort the orientation of fitted planes and bias the least-squares optimization process, thereby contaminating the estimation of parameters in other directions.
To address this, a decoupled 2D-3D estimation strategy is proposed to effectively prevent vertical inaccuracies from propagating into other components. Specifically, a pseudo-plane correspondence is introduced to replace the unreliable correspondence from DTM and provide an additional constraint in the GHM adjustment.
@@
Figure 6: Introduced pseudo-plane correspondence. $L_{i}$ and $M_{i}$ denote the LiDAR-derived and model-derived planes, respectively, and serve as facade-based correspondences for registration. The dashed lines represent the introduced pseudo-plane correspondence, which serves as a substitute for the original constraints derived from the LiDAR ground points and model ground points (DTM), indicated by red and blue round dots, respectively.
@@
$\mathbf{A}_{[n\times 7]}=\left.\frac{\partial f}{\partial\mathbf{X}}\right|_{\mathbf{X}_{0},\mathbf{V}_{0}},\qquad\mathbf{B}_{[n\times 3n]}=\left.\frac{\partial f}{\partial\mathbf{V}}\right|_{\mathbf{X}_{0},\mathbf{V}_{0}},\qquad\mathbf{C}_{[1\times 7]}=\left.\frac{\partial c}{\partial\mathbf{X}}\right|_{\mathbf{X}_{0}}.$ (11)
@@
The primary advantage of this strategy lies in its ability to perform reliable parameter estimation even in the absence of real ground-plane correspondences. More importantly, it effectively decouples the influence of low-quality observations–such as those derived from DTM data–from high-quality observations, preventing error propagation across components and maximizing the use of reliable facade-based correspondences.
On the one hand, such a design remains adaptable to ground data of varying types, accuracies, and acquisition sources. This adaptability becomes even more evident in scenarios with sloped or uneven ground surfaces, where simply constraining or omitting the parameter in a given direction–such as $t_{z}$ in the GHM–may still introduce additional errors. On the other hand, rather than omitting the estimation of $t_{z}$ and solving only for five parameters, the pseudo-plane strategy offers greater flexibility and adaptability, making it readily extendable to other facade orientations. Overall, the pseudo-plane strategy delivers a balanced design that preserves estimation robustness while maintaining adaptability to diverse scenarios.
@@ 3.4 Adaptable Vertical Translation Estimation
Figure 7: Vertical translation estimation.
In this way, the vertical translation $t_{z}$ can be easily estimated from DTM data while preserving a high degree of flexibility, benefiting from the aforementioned 2D-3D decoupled estimation strategy. For instance, if more accurate vertical information becomes available–such as GNSS observations, total station measurements, or other high-fidelity ground models–the DTM can be seamlessly substituted using the same method. This design greatly enhances the adaptability of the overall framework to heterogeneous data sources and ensures robustness under varying data availability conditions.
@@ 4.1 Data Preparation
To validate and evaluate the proposed L2M-Reg, four datasets from Munich (as shown in Figure 8), named as TUM0501 Building, Pinakothek, Street Building, and Restaurant, are collected, comprising LiDAR point clouds from TLS or MLS. Besides, the corresponding publicly accessible LoD2 models are obtained from the Bavarian Surveying Administration [^2] https://geodaten.bayern.de/opengeodata/. Additionally, one more dataset from another city, Ingolstadt, named Ingolstadt Store, is included in the experiment. These five datasets vary in point cloud size, location, building size, and data acquisition methods. It is worth noting that the two buildings, Restaurant and Ingolstadt Store, not only contain the wall plinths examined in this study, but also feature extensive glass elements on their exterior facades, including windows, street-facing glass doors, and glass display windows. In the Restaurant dataset, some of these glass surfaces are even covered with posters. All facades across the five datasets examined contain a certain degree of non-planar components, such as rainwater pipes, attached mailboxes, among others. As a result, the input point clouds used for plane extraction include not only the target planar regions but also several discrete non-planar structures.
Figure 8: Overview of five datasets. The upper part shows images of each building. LoD2 semantic models are in the middle, with planes colored for visual distinction. The lower part displays the corresponding point clouds, colored by intensity. For better visualization, the scale of each dataset in this figure is not directly comparable. Quantitative information is provided in Table 1.
In particular, due to the absence of ground data in the publicly accessible LoD2 models, ground information for the TUM0501 Building and Pinakothek datasets, which are part of the TUM2TWIN [^3] https://tum2t.win/ (Wysocki et al., 2026), is supplemented by publicly available DTM data with a grid width of 1 m from the Bavarian Surveying Administration. The Restaurant dataset also has the same DTM data. The Street Building and Ingolstadt Store datasets utilize additional georeferenced road models. Further details of the datasets are summarized in Table 1.
Table 1: Basic Information on Five Datasets
Number of Points Scanning Area Scanning System Ground Data Source Location
TUM0501 Building 2,272,810 $\approx$ 1,200 m2 Leica ScanStation P50 DTM Munich
Pinakothek 10,367,229 $\approx$ 3,600 m2 Leica ScanStation P50 DTM Munich
Street Building 5,000,509 $\approx$ 1,500 m2 Z+F FlexScan 22 Road Model Munich
Restaurant 24,116,910 $\approx$ 800 m2 Z+F FlexScan 22 DTM Munich
Ingolstadt Store 7,477,684 $\approx$ 60 m2 MoSES (Haigermoser et al., 2015) Road Model Ingolstadt
@@ 4.2 Experiment Design
In addition to the proposed L2M-Reg, five existing methods–GICP (Segal et al., 2009), TriICP (Chetverikov et al., 2002), KISS-ICP (Vizzo et al., 2023), PLADE (Chen et al., 2019), and Scantra (Wujanz et al., 2018)–are also applied to the five datasets for performance evaluation and comparison. GICP, TriICP, and KISS-ICP are chosen as leading robust ICP-variants, known for their reliability in fine-registration tasks. PLADE and Scantra are included as the two leading plane-based registration methods. Since these registration methods are inherently designed for point clouds rather than parametric models, the LoD2 models from the five datasets were converted into point clouds with uniform density to facilitate registration using the five methods mentioned above.
L2M-Reg was implemented in C++ using Point Cloud Library (PCL) (Rusu and Cousins, 2011). Standard implementations of GICP and TriICP in PCL (version 1.13.0) were utilized, while PLADE was implemented based on its open-source release [^4] https://github.com/chsl/PLADE (Chen et al., 2019). Scantra was tested on the basis of the latest version 3.4 released in 2025555https://www.technet-gmbh.com/en/products/scantra/scantra-news/scantra-release-34-static-kinematic-and-polar-are-ready-for-download/. KISS-ICP was reproduced using its open-source implementation  [^6] https://github.com/PRBonn/kiss-icp. All methods were executed under identical hardware conditions (Intel Xeon(R) W-2223 CPU@3.60GHz with 4 Cores and 64 GB of RAM).
@@ 4.3 Evaluation Metrics
where $n_{H}$ and $n_{V}$ denote the number of check points used for assessing horizontal and vertical error, respectively. For each check point $p$, $\delta_{\mathrm{M3C2}}(p)$ represents the M3C2 distance between $p$ and the reference model. Similarly, based on Eq. (14) and Eq. (15), the standard deviation of M3C2 distances in the horizontal and vertical directions can be computed by
@@ 4.4 Comparison Results
It is important to note that, unlike conventional evaluations based on the overall geometric closeness of the entire building, this study–motivated by the focus on model uncertainty–adopts the building footprint area as the primary basis for evaluation. This choice stems from the fact that, during LoD2 model generation, the near-ground footprint area is typically the most suited and accurately reflects the building’s true location, especially when there is a horizontal offset between plinth and facade. Accordingly, all subsequent evaluations in this study assess horizontal accuracy merely based on the distance between the aligned LiDAR point clouds and LoD2 models in the footprint area.
@@ 4.4.1 Qualitative Comparison
Figures 9, 10, 11, 12, and 13 show the registration results for the TUM0501, Pinakothek, Street Building, Restaurant, and Ingolstadt Store datasets, respectively. During testing, KISS-ICP, TriICP, GICP, and L2M-Reg consistently produced valid results for all datasets, whereas PLADE failed to do so for the TUM0501 Building, Restaurant, and Ingolstadt Store. Scantra exhibited suboptimal performance across all datasets. Possible reasons for these shortcomings are discussed in Section 5.
In addition, due to the high similarity in visual detail between KISS-ICP and other effective methods, visual distinction is very difficult. To improve clarity in visualization, results of KISS-ICP are not presented in the qualitative comparison of Figures 9 $\sim$ 13. The complete results of all methods can be seen in the subsequent quantitative evaluation.
@@
As shown in Figure 9, the TUM0501 Building dataset yields registration results using TriICP, GICP, and L2M-Reg. The results of PLADE are not included, as it can not produce valid results. In the horizontal direction, particularly within the building footprint areas outlined by dashed circles, L2M-Reg achieves the highest consistency compared to TriICP and GICP. In the vertical direction, it also shows superior alignment with the ground model, further highlighting its advantage in registration accuracy.
Figure 9: Registration performance of different methods in the TUM0501 Building dataset.
@@
Figure 10 shows registration results for the Pinakothek dataset. In the horizontal direction, L2M-Reg consistently demonstrates higher accuracy within the footprint area, as illustrated in Patch 1 to Patch 5 (outlined by dashed circles). Notably, in Patch 6, L2M-Reg exhibits a visible deviation from the reference model compared to the two ICP-based methods. This discrepancy is primarily attributed to scale inaccuracies in the LoD2 model, resulting in dimensional mismatches between the model and the point cloud. Despite this, L2M-Reg still achieves the globally optimal registration, as evidenced by Patch 1 being parallel to Patch 6, where L2M-Reg again provides a better global fit. In the vertical direction, as shown in Patch 7 to Patch 9, L2M-Reg demonstrates the best alignment.
Figure 10: Registration performance of different methods in the Pinakothek dataset.
@@
In the Street Building dataset, shown in Figure 11, L2M-Reg consistently delivers the most accurate alignment in the horizontal direction, particularly in Patch 1 and Patch 4. Similar slight dimensional mismatches between the models and the point clouds are observed in Patch 2 and Patch 3, as seen in the Pinakothek dataset. These mismatches are again attributed to scale inaccuracies introduced in the original models. Nevertheless, L2M-Reg achieves a globally optimal result, with the registered point clouds (in red) centrally aligned between the two corresponding facades of the LoD2 models in Patch 2 and Patch 3.
@@
In the vertical direction (e.g., Patch 6), L2M-Reg also achieves the best alignment in most regions. However, in certain areas like Patch 5, it does not produce optimal results. This results from the road being originally described according to the OpenDRIVE standard (ASAM, 2024; Kutsch et al., 2024), which models individual lanes with overlapping surface geometries and continuous curvature. The parametric surface representations are then sampled to explicit planar surface geometries as part of the conversion to CityGML 3.0 (Schwab et al., 2020). Although unintended overlapping surface geometries introduce minor instability in the evaluation of vertical accuracy, this demonstrates that L2M-Reg can also directly incorporate further object classes of the model, such as the road surface, into the vertical alignment process. Since CityGML 3.0 enables a comprehensive and redundancy-free representation of the road space, such ambiguities are not introduced in native mapping processes.
As shown in Figure 12, the Restaurant dataset yields registration results using TriICP, GICP, and L2M-Reg. PLADE is not able to give valid results in this case. In the horizontal direction, particularly within the building footprint areas outlined by dashed circles in Patch 1 to Patch 3, L2M-Reg achieves the highest consistency. In the vertical direction, it also shows superior alignment, as illustrated in Patch 5. However, in Patch 4, its performance falls slightly short of the optimal method.
Figure 12: Registration performance of different methods in the Restaurant dataset.
Figure 13 shows the registration results for the Ingolstadt Store dataset. In the horizontal direction, L2M-Reg consistently demonstrates the best alignment within the footprint area, as illustrated in Patch 1 to Patch 4 (outlined by dashed circles). In the vertical direction, as shown in Patch 5 to Patch 6, L2M-Reg also demonstrates the best alignment between point clouds and road model.
Figure 13: Registration performance of different methods in the Ingolstadt Store dataset.
In general, the qualitative comparison results indicate that L2M-Reg outperforms the other methods in both horizontal and vertical consistency in all datasets. More detailed quantitative evaluation results are presented in Section 4.4.2.